# Unifying Summary Statistic Selection for Approximate Bayesian Computation

### v3

@@ 4.6 Model selection
Prangle et al. (2014) used logistic regression to learn summaries that can discriminate between different models: The predicted class probabilities. Similarly, Merten et al. (2019) applied deep convolutional neural networks to weak lensing maps to learn features that can discriminate between nine different cosmological models, although not in the context of ABC. Such probabilistic approaches to model classification are equivalent to minimizing the EPE: Consider a one-hot encoding of the model index such that $\theta_{j}=1$ if model $j$ generated the data and $\theta_{j}=0$ otherwise. The log posterior is thus
@@
where $r$ is the number of models under consideration, and $f\left(\theta_{j}=1\mid t\right)$ is the posterior probability that the data were generated by model $j$. Equation 12 is familiar as the negative cross-entropy loss widely used for multiclass classification in machine learning (Bishop, 2006, Ch. 4). In other words, any machine learning classifier that minimizes the cross-entropy loss for model selection minimizes the EPE of the model labels.
@@ 4.7 Conditional posterior density estimation
As briefly discussed in Section 3, recent approaches to likelihood-free inference based on conditional density estimation minimize the EPE using the mini-batch estimator in Eq. 4 as a loss function (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Radev et al., 2022). These methods are appealing because they can automatically compress large datasets although at the cost of having to choose an architecture for the density estimator which is an active area of research (Papamakarios et al., 2021). Chan et al. (2018) proposed exchangeable neural networks such that the output is invariant under permutations of i.i.d. data generated by the model. While neural density estimators can in principle learn such invariances, it is beneficial to encode symmetries in the architecture to improve efficiency and reduce the amount of training data required. In practice, conditional density estimators can have computational advantages over rejection ABC because they interpolate the posterior density in the parameter space, requiring fewer simulations (Papamakarios & Murray, 2016). However, they cannot offer the same asymptotic guarantees as ABC: As the tolerance parameter of the acceptance kernel is reduced, the sampling distribution converges to the true posterior (Beaumont, 2019).
@@ 4.8 Partial least squares regression
Wegmann et al. (2009) obtained summaries using partial least squares regression (PLSR), a latent variable model for supervised dimensionality reduction. The method projects data $z$ (or candidate summaries) to a latent space such that the embeddings are most predictive of the parameters $\theta$ as measured by the $L_{2}$ norm. Instead of the predictions of the model as in Section 4.4, the latent variables are used as summaries. The number of latent components is chosen using leave-one-out cross-validation based on the ability of the model to predict parameters. Similar to the subset selection methods discussed in Sections 4.1 and 4.2, the number of components chosen by cross-validation is determined by computational constraints: For sufficiently large reference tables, the dimensionality of the candidate summaries is maintained provided each candidate summary encodes some information, however weak.
@@ 5.1 Evaluation criteria and model architecture for nonlinear methods
Figure 4: Mixture density networks with a bottleneck can learn informative summaries. The left section illustrates the training procedure: $p$-dimensional parameters $\theta$ and synthetic data $z$ are drawn from the prior $\pi$ and simulator $g$, respectively. Synthetic data are compressed to summaries using a neural compressor $t$, and a mixture density network (MDN) $h$ estimates a posterior approximation $\hat{f}\left(\theta\mid t(z)\right)$, where $\mathcal{F}$ are the supported posteriors, e.g. MDNs with certain component distributions. The network is trained by minimizing the negative log probability (NLP) loss. The right section illustrates approximate Bayesian computation using learned summaries: The trained compressor evaluates summaries of observed data $y$, and parameter samples are accepted if corresponding simulated summaries $t\left(z\right)$ are sufficiently close to observed summaries $t\left(y\right)$. The table lists the type of data $\mathbb{D}$ and compressor architecture for each experiment (see Sections 5.2, 5.3 and 5.4 for details).
@@
Summaries minimizing the EPE are appealing, but a concrete algorithm is required to make them useful in practice. We employed a conditional mixture density network (MDN) (Papamakarios & Murray, 2016) with a bottleneck akin to an autoencoder (Kramer, 1991). The network comprises two parts: First, for fair comparison, we used the same neural compressor as for nonlinear posterior mean estimation. Consequently, the number of summaries $q$ is equal to the number of parameters $p$, although, in general, a different number of summaries $q>p$ could be chosen (Chen et al., 2021). Second, we extended the network with a conditional MDN to estimate the posterior density given only the summaries. The composite network was trained by minimizing the Monte Carlo estimate of the EPE defined in Eq. 4. After training, the bottleneck architecture ensures any information that may be useful for minimizing the EPE is captured by the output of the compressor; we dub this approach MDN compression. The inference pipeline is illustrated in Fig. 4. A similar approach was used by Jeffrey et al. (2020) for summaries fed to a likelihood estimation network. Radev et al. (2022) used a similar architecture of compression and density estimation networks, although using a normalizing flow for the latter. They used 128 summaries which is prohibitively large for ABC.
@@
For each experiment, we generated a training, validation, and test set by sampling from the prior predictive distribution. Neural compressors were trained by mini-batch gradient descent using the Adam optimizer with default parameters and an initial learning rate of $10^{-2}$ (Kingma & Ba, 2015). The learning rate was decreased by an order of magnitude if the loss evaluated on the validation set did not decrease for ten consecutive epochs; training was stopped if it did not decrease for twenty consecutive epochs.
@@
After extracting summaries for each example in the test set, we obtained samples from the approximate posterior $\tilde{f}\left(\theta\mid t\left(y\right)\right)$ in two steps: First, we standardized summaries to ensure a common scale and evaluated the Euclidean distance $d\left(t\left(y\right),t\left(z_{i}\right)\right)$ between each example $y$ and the $i^{\text{th}}$ element of the training set $z_{i}$. Second, we accepted a small fraction of the reference table as posterior samples such that they had the smallest distance to each example (Beaumont, 2019), i.e. the training set served as the reference table. In addition to ABC, we drew samples from the prior as a baseline as well as directly from the MDNs trained to obtain EPE-minimizing summaries.
@@
where $\tilde{\theta}_{i}$ denotes the $i^{\text{th}}$ sample from the approximate posterior and $s$ is the number samples. This metric has been widely used in the ABC literature to evaluate summary extraction methods (Joyce & Marjoram, 2008; Nunes & Balding, 2010; Fearnhead & Prangle, 2012; Blum et al., 2013; Burr & Skurikhin, 2013; Jiang et al., 2017). It measures how concentrated approximate posterior samples are around the true parameter value $\theta$ (Bishop, 2006, Ch. 3). The RMISE is a suitable metric for unimodal but not multimodal posteriors, as illustrated in panel (b) of Fig. 2. Second, to address this shortcoming, we also evaluated the negative log probability (NLP) using Gaussian kernel density estimation with bandwidth chosen by Scott’s rule (Scott, 2015).
@@ 5.2 Benchmark model
Figure 5: A conditional mixture density network (MDN) that minimizes the expected posterior entropy learns highly informative summaries. Panel (a) shows the likelihood for a particular parameter value $\theta_{0}\approx 1.6$ together with a rug plot for $n=10$ samples $y_{0}$. Panel (b) shows the exact posterior $f\left(\theta\mid y\right)$ given the data in (a) together with the learned mixture density. While the two-component mixture is not flexible enough to approximate the posterior well, it learns highly informative summaries: MDN-compressed ABC samples using these summaries are shown as a histogram. Panel (c) shows the learned summary $t(\cdot)$; markers correspond to the data in (a). The dashed line is a fit to the learned summaries based on the first three even moments of $y$. Panel (d) illustrates the relationship between the mixture density approximation and the summary as a heat map; lighter colours indicated higher posterior density.
@@
We considered a benchmark model with multimodal posterior set up to be challenging for extracting summaries. The model has a tractable likelihood that allowed us to compare the approximate posterior given summaries with exact posterior samples. In particular, we sampled a univariate parameter $\theta$ from the standard normal distribution and drew $n=10$ independent samples from the mixture distribution
@@
$y_{i}\mid\theta\sim\frac{1}{2}\sum_{u\in\left\{-1,1\right\}}\mathsf{Normal}\left(u\times\tanh\theta,1-\tanh^{2}\theta\right),$
@@
as illustrated in panel (a) of Fig. 5. We also sampled a standard normal distractor (uninformative noise) for each observation $i$ such that the full dataset $Y$ is a matrix with $n$ rows and $2$ columns. Learning or selecting summaries is non-trivial because all elements of $Y$ have zero mean and unit variance under the generative model irrespective of the parameter $\theta$. The first moment is zero by symmetry; the second moment of each mixture component is $\mathbb{E}\left[y^{2}\right]=\mathbb{E}\left[y\right]^{2}+\operatorname{var}y=\tanh^{2}\theta+1-\tanh^{2}\theta=1$ such that the mixture has unit variance. Sampling from the prior predictive distribution, we generated training, validation, and test sets of $10^{6}$, $10^{4}$, and $10^{3}$ independent realizations, respectively. The test set was used to evaluate and compare different methods. We employed the likelihood-based inference framework Stan (Carpenter et al., 2017) to draw $1{,}000$ posterior samples for each example in the test set (see App. C for details). These samples formed the “ground truth” which we compared other methods to.
@@
For ABC using candidate summaries, the CPE minimization method (Nunes & Balding, 2010), PLS (Wegmann et al., 2009), and linear posterior mean estimation (Fearnhead & Prangle, 2012), we used the first three even moments of each column of $Y$ as candidate summaries, giving rise to six statistics in total. Odd moments are not informative as the likelihood is symmetric, and we did not include them in our set of candidate summaries.
@@
For the nonlinear posterior mean approximation (Jiang et al., 2017), we used a multilayer perceptron (MLP) that acts on each row of $Y$ independently before compressing to a scalar summary. This architecture shares weights across all observations and is permutation invariant (Chan et al., 2018). The MLP has three fully connected layers (16, 16, and 1 hidden units) followed by mean-pooling across the $n$ observations; we used $\tanh$ activation functions throughout. The network was implemented in PyTorch (Paszke et al., 2019) and trained as described in the preceding section with a mini-batch size of $512$.
@@
Finally, we used a conditional MDN with $k=2$ Gaussian components to estimate the posterior and learn MDN-compressed summaries (Bishop, 1994; Papamakarios & Murray, 2016). To evaluate mixture logits $\eta(t)$, locations $\mu(t)$, and log-scales $\kappa(t)$ as a function of the summary $t$ we used independent two-layer MLPs (16 and $k$ hidden layers). The approximate posterior is thus
@@
Figure 6: The quality of summaries significantly impacts the fidelity of posteriors. Panels (a), (b), and (c) report the negative log probability loss and root mean integrated squared error of different methods for the benchmark, coalescent, and growing tree experiments, respectively. Error bars are standard errors based on a test set of 1,000 i.i.d. samples for each experiment. Expert summaries and PLS perform poorly for the growing tree experiment and are indicated as off-the-chart by arrows. See Sections 5.2, 5.3 and 5.4 for details on the expert summaries for the benchmark, coalescent, and growing tree models, respectively.
@@
A comparison of the performance of different methods based on $1{,}000$ approximate posterior samples ($0.1\%$ of the reference table) is shown in panel (a) of Fig. 6 (see Appendix C in the appendix for a table of results). We report the RMISE for completeness, but it is a poor metric for multimodal posteriors. For example, a point mass at $\theta=0$ would have $\text{RMISE}=1$, lower than any of the methods we considered. As expected, linear and nonlinear posterior mean estimators performed worst in terms of NLP because the posterior is bimodal. Because of its flexibility, the nonlinear estimator was able to accurately estimate the posterior mean $\mathbb{E}_{\theta\sim f\left(\theta\mid z\right)}\left[\theta\right]=0$ which, ironically, led to the worst performance: The NLP is equal to the prior entropy ($1.42$). The linear estimator performed better because the regression coefficients are entirely determined by noise in the training set, i.e. the scalar summary is a random projection of the candidate summaries. Similarly, extracting features using PLS regression is driven by noise: Here, three random orthogonal projections of candidate features were selected based on five-fold cross-validation, allowing PLS to outperform both linear and nonlinear regression. Minimizing the conditional posterior entropy and using candidate summaries without selection performed similarly and better than regression-based approaches.
@@
MDN-compressed ABC performed almost as well as the gold standard likelihood-based inference and better than samples drawn directly from the MDN, as illustrated in panel (b) of Fig. 5 for a particular example. The architecture of the MDN is too restrictive to approximate the posterior well. Increasing the number of components to $k=10$ provides a better approximation ($\text{NLP}=1.05\pm 0.01$) with the same performance as both the likelihood-based approach and MDN-compressed ABC. Here, we deliberately restricted the architecture to illustrate that ABC with good summaries can remain competitive because it does not rely on parametric assumptions about the density.
@@
Due to the simplicity of the benchmark problem, we can inspect the MDN and learned summary, as shown in panel (c). The appropriate summary is obvious in retrospect: It should discriminate between data $y$ clustered around ±$\pm 1$ (corresponding to large absolute values of $\theta$) and data near zero or large absolute value (corresponding to small absolute values of $\theta$). A linear fit to $t(y)$ given candidate summaries illustrates that they are rich enough to provide a high-quality summary in principle, but most methods struggled to extract the information. We obtained the fit by minimizing squared residuals on the interval $\left(-3,3\right)$ weighted by the prior density. Finally, the density of the MDN, shown in panel (d), exhibits the expected behaviour: Large summaries give rise to unimodal distributions centred at the origin, and small summaries yield bimodal posterior estimates.
@@
The choice of compressor architecture is not unique. For example, we could have included further layers after the mean-pooling operation or used a fully-connected network throughout. However, using the mean-pooled latent features has a unique advantage: They are unbiased estimates of the population mean of the features independent of sample size. The architecture was motivated by the observation that the likelihood of exponential family distributions can be expressed in terms of sums (or means) of transformations of the data. We thus expect the learned summaries to remain informative for different sample sizes. To test this hypothesis, we repeated the analysis with $n=100$ instead of $n=10$ observations per example. The NLPs are smaller because we had access to more data: $0.68\pm 0.01$ for likelihood-based inference and $0.70\pm 0.01$ for MDN-compressed ABC, where the MDN was trained on the larger dataset using the same methodology as before. Running MDN-compressed ABC with the network trained on the smaller dataset yielded a NLP of $0.72\pm 0.01$, i.e. the performance is almost indistinguishable from the network trained on the larger dataset despite being exposed to an order of magnitude fewer observations. Importantly, the MDN itself cannot achieve this generalization because the model was trained on data with a fixed sample size. BayesFlow seeks to provide amortized inference even for variable sample sizes although at the cost of further simulations (Radev et al., 2022).

### v4

@@ 4.6 Model selection
Prangle2014 used logistic regression to learn summaries that can discriminate between different models: The predicted class probabilities. Similarly, Merten2019 applied deep convolutional neural networks to weak lensing maps to learn features that can discriminate between nine different cosmological models, although not in the context of ABC. Such probabilistic approaches to model classification are equivalent to minimizing the EPE: Consider a one-hot encoding of the model index such that $\theta_{j}=1$ if model $j$ generated the data and $\theta_{j}=0$ otherwise. The log summary posterior is thus
@@
where $r$ is the number of models under consideration, and $f\left(\theta_{j}=1\mid t\right)$ is the posterior probability that the data were generated by model $j$. Equation 13 is familiar as the negative cross-entropy loss widely used for multiclass classification in machine learning [Bishop2006, Ch. 4]. In other words, any machine learning classifier that minimizes the cross-entropy loss for model selection minimizes the EPE of the model labels.
@@ 4.7 Conditional posterior density estimation
As briefly discussed in Section 3, recent approaches to likelihood-free inference based on conditional density estimation minimize the EPE using the mini-batch estimator in Eq. 5 as a loss function [Papamakarios2016, Lueckmann2017, Radev2022BayesFlow]. These methods are appealing because they can automatically compress large datasets although at the cost of having to choose an architecture for the density estimator which is an active area of research [Papamakarios2021]. Chan2018 proposed exchangeable neural networks such that the output is invariant under permutations of i.i.d. data generated by the model. While neural density estimators can in principle learn such invariances, it is beneficial to encode symmetries in the architecture to improve efficiency and reduce the amount of training data required. In practice, conditional density estimators can have computational advantages over rejection ABC because they interpolate the posterior density in the parameter space, requiring fewer simulations [Papamakarios2016]. However, they cannot offer the same asymptotic guarantees as ABC: As the tolerance parameter of the acceptance kernel is reduced, the sampling distribution converges to the summary posterior [Beaumont2019].
@@ 4.8 Partial least squares regression
Wegmann2009PartialLeastSquares obtained summaries using partial least squares regression (PLSR), a latent variable model for supervised dimensionality reduction. The method projects data $z$ (or candidate summaries) to a latent space such that the embeddings are most predictive of the parameters $\theta$ as measured by the $L_{2}$ norm. Instead of the predictions of the model as in Section 4.4, the latent variables are used as summaries. The number of latent components is chosen using leave-one-out cross-validation based on the ability of the model to predict parameters. Similar to the subset selection methods discussed in Sections 4.1 and 4.2, the number of components chosen by cross-validation is determined by computational constraints: For sufficiently large reference tables, the dimensionality of the candidate summaries is maintained provided each candidate summary encodes some information, however weak.
@@ 5.1 Evaluation criteria and model architecture for nonlinear methods
Figure 4: Mixture density networks with a bottleneck can learn informative summaries. The stack left of the compressor $t$ illustrates the training data generation and MDN training procedure: $p$-dimensional parameters $\theta$ and synthetic data $z$ are drawn from the prior $\pi$ and simulator $g$, respectively. Synthetic data are compressed to summaries using a compressor $t$. The stack right of the compressor $t$ illustrates approximate Bayesian computation using learned summaries: The compressor evaluates summaries of observed data $y$, and parameter samples are accepted if corresponding simulated summaries $t\left(z\right)$ are sufficiently close to observed summaries $t\left(y\right)$. The red dashed box indicates components specific to training MDN compression: A mixture density network (MDN) $h$ estimates a posterior approximation $\hat{f}\left(\theta\mid t(z)\right)$ given learned summaries $t(z)$. Here, $\mathcal{F}$ are the supported posteriors, e.g. MDNs with certain component distributions. The network is trained by minimizing the negative log probability (NLP) loss. The table lists the type of data $\mathbb{D}$ and compressor architecture for each experiment (see Sections 5.2, 5.3 and 5.4 for details).
@@
Summaries minimizing the EPE are appealing, but a concrete algorithm is required to make them useful in practice. We employed a conditional mixture density network (MDN) [Papamakarios2016] with a bottleneck akin to an autoencoder [Kramer1991Autoencoder]. The network comprises two parts: First, for fair comparison, we used the same neural compressor as for nonlinear posterior mean estimation. Consequently, the number of summaries $q$ is equal to the number of parameters $p$, although, in general, a different number of summaries $q>p$ could be chosen [Chen2021MutualInfo]. Second, we extended the network with a conditional MDN to estimate the posterior density given only the summaries. The whole network comprising compressor and MDN was trained end-to-end by minimizing the Monte Carlo estimate of the EPE defined in Eq. 5. After training, the bottleneck architecture ensures any information that may be useful for minimizing the EPE is captured by the output of the compressor; we dub this approach MDN compression. The inference pipeline for all methods is illustrated in Fig. 4. For MDN compression, the compressor is trained by jointly optimizing a mixture density network on simulated data (shown in red) to minimize the expected posterior entropy. Once trained, summaries are extracted and used in ABC like other methods. A similar approach was used by Jeffrey2020NeuralSummaries for summaries fed to a likelihood estimation network. Radev2022BayesFlow used a similar architecture of compression and density estimation networks, although using a normalizing flow for the latter. They used 128 summaries which is prohibitively large for ABC.
@@
For each experiment, we generated a training, validation, and test set by sampling from the prior predictive distribution. Neural compressors were trained by mini-batch gradient descent using the Adam optimizer with default parameters and an initial learning rate of $10^{-2}$ [Kingma2015]. The learning rate was decreased by an order of magnitude if the loss evaluated on the validation set did not decrease for ten consecutive epochs; training was stopped if it did not decrease for twenty consecutive epochs.
@@
After extracting summaries for each example in the test set, we obtained samples from the approximate posterior $\tilde{f}\left(\theta\mid t\left(y\right)\right)$ in three steps: First, to ensure a common scale across summaries, we standardized them independently to have zero mean and unit variance based on the training set. Second, we evaluated the Euclidean distance $d_{i}$ between standardized summaries of each example $y$ and the $i^{\text{th}}$ element of the training set $z_{i}$. Third, we accepted a small fraction of the training set as posterior samples such that they had the smallest distance to each example [Beaumont2019], i.e. the training set served as the reference table. The same training, validation and test sets were used for different methods for fair comparison. In addition to ABC, we drew samples from the prior as a baseline as well as directly from the MDNs trained to obtain EPE-minimizing summaries.
@@
where $\tilde{\theta}_{i}$ denotes the $i^{\text{th}}$ sample from the ABC posterior and $s$ is the number of samples. This metric has been widely used in the ABC literature to evaluate summary extraction methods [Joyce2008, Nunes2010, Fearnhead2012, Blum2013, Burr2013Calibration, Jiang2017]. It measures how concentrated ABC posterior samples are around the true parameter value $\theta$ [Bishop2006, Ch. 3]. The RMISE is a suitable metric for unimodal but not multimodal posteriors, as illustrated in panel (b) of Fig. 2. Second, to address this shortcoming, we also evaluated the negative log probability (NLP) using kernel density estimation. Specifically,
$\text{NLP}=-\log\left[\frac{1}{s}\sum_{i=1}^{s}K_{h}\left(\tilde{\theta}_{i}-\theta\right)\right],$
where $K_{h}$ is a Gaussian kernel with bandwidth $h$ chosen by Scott’s rule [Scott2015MultivariateDensityEstimation]. For each experiment, metrics reported in Fig. 6 and Table 2 in the appendix are averaged over the corresponding test set.
@@ 5.2 Benchmark model
Figure 5: A conditional mixture density network (MDN) that minimizes the expected posterior entropy learns highly informative summaries. Panel (a) shows the likelihood for the true parameter $\theta^{*}\approx 1.6$ that generated the example dataset $y$ together with a rug plot for the $n=10$ observations $y_{\bullet 1}$. Panel (b) shows the true posterior $f\left(\theta\mid y\right)$ together with the learned posterior density estimator. While the two-component mixture is not flexible enough to approximate the true posterior well, it learns highly informative summaries: MDN-compressed ABC samples using these summaries are shown as a histogram. Panel (c) shows the learned summary function $t:\mathbb{R}^{10\times 2}\to\mathbb{R}$ which maps the full data matrix to a scalar; the plot shows $t(y)$ as a function of the first column values $y_{\bullet 1}$ (the informative data, with the second column being uninformative noise). The dashed line shows how $t$ can be approximated using polynomial basis functions of the candidate summaries (the first three even moments). Panel (d) illustrates the relationship between the posterior density estimator and the summary as a heat map; lighter colours indicate higher posterior density.
@@
We considered a benchmark model with multimodal true posterior set up to be challenging for extracting summaries. The model has a tractable likelihood that allowed us to compare the posterior given summaries with true posterior samples. In particular, we sampled a univariate parameter $\theta$ from the standard normal distribution and drew $n=10$ independent samples from the mixture distribution
@@
$z_{i1}\mid\theta\sim\frac{1}{2}\sum_{u\in\left\{-1,1\right\}}\mathsf{Normal}\left(u\times\tanh\theta,1-\tanh^{2}\theta\right),$
@@
as illustrated in panel (a) of Fig. 5. We also sampled a standard normal distractor $z_{i2}$ (uninformative noise) for each observation $i$ such that the full dataset $z=[z_{ij}]$ is a matrix with $n$ rows and $2$ columns. Learning or selecting summaries is non-trivial because all elements of $z$ have zero mean and unit variance under the generative model irrespective of the parameter $\theta$. The first moment is zero by symmetry; the second moment of each mixture component is $\mathbb{E}\left[z_{i1}^{2}\right]=\mathbb{E}\left[z_{i1}\right]^{2}+\operatorname{var}z_{i1}=\tanh^{2}\theta+1-\tanh^{2}\theta=1$ such that the mixture has unit variance. Sampling from the prior predictive distribution, we generated training, validation, and test sets of $10^{6}$, $10^{4}$, and $10^{3}$ independent realizations, respectively. The test set was used to evaluate and compare different methods. We employed the likelihood-based inference framework Stan [Carpenter2017] to draw 1,0001{,}000 posterior samples for each example in the test set (see App. C for details). These samples formed the gold standard which we compared other methods to. Figure 5 illustrates the learned summaries for a particular example dataset $y$ generated with true parameter $\theta^{*}\approx 1.6$.
@@
For ABC using candidate summaries, the CPE minimization method [Nunes2010], PLS [Wegmann2009PartialLeastSquares], and linear posterior mean estimation [Fearnhead2012], we used the first three even moments of each column of $z$ as candidate summaries, giving rise to six statistics in total. Odd moments are not informative as the likelihood is symmetric, and we did not include them in our set of candidate summaries.
@@
For the nonlinear posterior mean approximation [Jiang2017], we used a multilayer perceptron (MLP) that acts on each row of $z$ independently before compressing to a scalar summary. This architecture shares weights across all observations and is permutation invariant [Chan2018]. The MLP has three fully connected layers (16, 16, and 1 hidden units) followed by mean-pooling across the $n$ observations; we used $\tanh$ activation functions throughout. The network was implemented in PyTorch [Paszke2019] and trained as described in the preceding section with a mini-batch size of $512$.
@@
Finally, we used a conditional MDN with $k=2$ Gaussian components to estimate the posterior and learn MDN-compressed summaries [Bishop1994, Papamakarios2016]. To evaluate mixture logits $\eta(t)$, locations $\mu(t)$, and log-scales $\kappa(t)$ as a function of the summary $t$ we used independent two-layer MLPs (16 and $k$ hidden layers). The posterior density estimator is thus
@@
Figure 6: The quality of summaries significantly impacts the fidelity of posteriors. Panels (a), (b), and (c) report the negative log probability loss and root mean integrated squared error of different methods for the benchmark, coalescent, and growing tree experiments, respectively. “MDN” refers to drawing samples directly from the learned mixture density network, “likelihood-based” refers to samples obtained with Stan for the benchmark problem and pseudo-marginal MCMC for the growing tree model, and all other methods use ABC after extracting summaries. Error bars are standard errors based on a test set of 1,000 i.i.d. samples for each experiment. Expert summaries and PLS perform poorly for the growing tree experiment and are indicated as off-the-chart by arrows. See Sections 5.2, 5.3 and 5.4 for details on the expert summaries for the benchmark, coalescent, and growing tree models, respectively.
@@
A comparison of the performance of different methods based on 1,0001{,}000 approximate posterior samples ($0.1\%$ of the reference table) is shown in panel (a) of Fig. 6 (see Table 2 in the appendix for a table of results). We report the RMISE for completeness, but it is a poor metric for multimodal posteriors. For example, a point mass at $\theta=0$ would have $\text{RMISE}=1$—lower than any of the methods we considered. As expected, linear and nonlinear posterior mean estimators performed worst in terms of NLP because the posterior is bimodal. Because of its flexibility, the nonlinear estimator was able to accurately estimate the posterior mean $\mathbb{E}_{\theta\sim f\left(\theta\mid z\right)}\left[\theta\right]=0$ which, ironically, led to the worst performance: The NLP is equal to the prior entropy ($1.42$). The linear estimator performed better because the regression coefficients are entirely determined by noise in the training set, i.e. the scalar summary is a random projection of the candidate summaries. Similarly, extracting features using PLS regression is driven by noise: Here, three random orthogonal projections of candidate features were selected based on five-fold cross-validation, allowing PLS to outperform both linear and nonlinear regression. Minimizing the conditional posterior entropy and using candidate summaries without selection performed similarly and better than regression-based approaches.
@@
MDN-compressed ABC performed as well as the gold standard likelihood-based inference ($\text{NLP}=1.05\pm 0.01$) and better than samples drawn directly from the MDN ($\text{NLP}=1.08\pm 0.02$), as illustrated in panel (b) of Fig. 5 for a particular example. While the bottleneck forces the network to compress data to an informative summary statistic, the architecture of the MDN with only two mixture components is too restrictive to approximate the true posterior well. Increasing the number of components to $k=10$ provides a better approximation with the same performance as both the likelihood-based approach and MDN-compressed ABC. Here, we deliberately restricted the architecture to illustrate that ABC with good summaries can remain competitive because it does not rely on parametric assumptions about the density.
@@
Due to the simplicity of the benchmark problem, we can inspect the MDN and learned summary, as shown in panel (c). The appropriate summary is obvious in retrospect: It should discriminate between data $z_{\bullet 1}$ clustered around $\pm 1$ (corresponding to large absolute values of $\theta$) and data near zero or large absolute value (corresponding to small absolute values of $\theta$). The dashed line shows a polynomial approximation of the learned summary $t$ using the candidate summaries (the first three even moments of each column) as basis functions. This fit illustrates that the candidate summaries are rich enough to provide a high-quality summary in principle, but most methods struggled to extract the information. We obtained the fit by minimizing squared residuals on the interval $\left(-3,3\right)$ weighted by the prior density. Finally, the density of the MDN, shown in panel (d), exhibits the expected behaviour: Large summaries give rise to unimodal distributions centred at the origin, and small summaries yield bimodal posterior estimates.
@@
The choice of compressor architecture is not unique. For example, we could have included further layers after the mean-pooling operation or used a fully-connected network throughout. However, using the mean-pooled latent features has several advantages: First, the number of compressor parameters is independent of the sample size. Second, they are unbiased estimates of the population mean of the features independent of sample size. The architecture was motivated by the observation that the likelihood of exponential family distributions can be expressed in terms of sums (or means) of transformations of the data and preserves the i.i.d. structure required to connect Fisher information maximization with EPE minimization as discussed in Section 4.3. We thus expect the learned summaries to remain informative for different sample sizes. To test this hypothesis, we repeated the analysis with $n=100$ instead of $n=10$ observations per example. The NLPs are smaller because we had access to more data: $0.68\pm 0.01$ for likelihood-based inference and $0.70\pm 0.01$ for MDN-compressed ABC, where the MDN was trained on the larger dataset using the same methodology as before. Running MDN-compressed ABC with the network trained on the smaller dataset yielded a NLP of $0.72\pm 0.01$, i.e. the performance is almost indistinguishable from the network trained on the larger dataset despite being exposed to an order of magnitude fewer observations. Importantly, the posterior density estimator itself cannot achieve this generalization because the model was trained on data with a fixed sample size. BayesFlow seeks to provide amortized inference even for variable sample sizes although at the cost of further simulations [Radev2022BayesFlow].